OpenStudy (meilendurcer):

A theater has 500 seats. Three-fourths of the seats are filled. How many seats are filled? i have 384

OpenStudy (nubeer):

number of seats filled = (3/4) * 500

OpenStudy (anonymous):

75% of 500 -> 0.75 * 500 = (3/4) * 500 = 375

OpenStudy (anonymous):

one fourth of 500 is 125 three times 125 is your answer